How do I solve this?
The line through (3, 4) that is parallel to the line 5x + 3y = 9

Answer by mananth(16949)

The line through (3, 4) that is parallel to the line 5x + 3y = 9
5x+3y=9
re write the equation in slope intercept form
3y=-5x+9
y=(-5/3)x+3
so the slope m = -5/3
since the required line is parallel the slope will be the same.
Now find b by plugging th value (3,4) & (-5/3)
4=(-5/3)*3+b
4--5+b
b=9
So the equation will be
y = -5x/3 +9
....
m.ananth@hotmail.ca